# Analysis of a finite element method for second order uniformly elliptic PDEs in non-divergence form

### v2

@@ 3.1. Global $W^{2,p}$ estimate of linear elliptic PDE in non-divergence form with uniformly continuous AA
Remark 3.1.
Though it is inspired by that of [24, Lemma $9.16$], the proof of Lemma 3.1 avoids using the common technique of “flattening” $\partial\Omega$, which is critical in [24, Lemma $9.16$]. Thus Lemma 3.1 is valid on convex polyhedra in $\mathbb{R}^{3}$.
@@
Since $\boldsymbol{x}_{0}\in\overline{\Omega}$ is chosen arbitrarily, we can conclude that $\phi\in W^{2,q}(\Omega)$ where $q=\min(\dfrac{3p}{3-p},2)>\frac{3}{2}$. So $\phi\in C^{0}(\overline{\Omega})$. Furthermore, by the interior estimate in [24, Lemma $9.16$], we have $\phi\in W_{\text{loc}}^{2,3}(\Omega)$ since $A:D^{2}\phi+\boldsymbol{b}\cdot\nabla\phi+c\phi=0$ in $\Omega$. Therefore, by the ABP estimate [24, Theorem $9.1$], we can conclude that $\phi=0$ in $\Omega$. ∎
@@
When $d=2$, $W^{2,p}(\Omega)\Subset C^{0}(\overline{\Omega})$ due to Sobolev’s embedding theorem. Furthermore, by the interior estimate in [24, Lemma $9.16$], we have $u\in W_{\text{loc}}^{2,2}(\Omega)$. Therefore, Theorem 3.2 can be proven by mimicking the proofs of [34, Theorem $1.1$] and [34, Corollary $1.2$] with the global $W^{2,p}$ estimate of Poisson equation (see [7, Proposition $2.1$]) and the Alexandroff–Bakelman–Pucci estimate estimate [24, Theorem $9.1$].
Now we focus on the case $d=3$. Due to [34, Corollary $1.2$], Theorem 3.2 holds if we further assume $\Omega$ is a convex polyhedra in $\mathbb{R}^{3}$ and $\frac{3}{2}<p\leq 2$. In fact, if we replace [10, Corollary $3.12$] by Lemma 2.1 in the proof of [34, Theorem $1.1$], it is easy to have that for any $1<p\leq 2$ and for any $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
@@
As the proof of [34, Corollary $1.2$], Theorem 3.2 will be proven via (27) and the proof of contraction with the uniqueness of solution in $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$:
@@
When $\frac{3}{2}<p\leq 2$, it is straightforward by the Alexandroff–Bakelman–Pucci estimate estimate with the fact $W^{2,p}(\Omega)\Subset C^{0}(\overline{\Omega})$ (see the proof of [34, Corollary $1.2$]). Therefore, in order to complete the proof of Theorem 3.2, we only need the uniqueness for $1<p\leq\frac{3}{2}$, which is provided by Lemma 3.1. Therefore, we can conclude that the proof of Theorem 3.2 is complete. ∎
@@ 3.2. Stability analysis of FEM for linear elliptic PDE in non-divergence form with uniformly continuous AA
We define a linear operator $\mathcal{L}_{\tilde{A}_{0},h}:\rightarrow W_{h}^{1}\rightarrow V_{h}$ as
@@
(30) $\|w_{h}-w\|_{W^{1,p}(\Omega)}\leq C\sup_{\chi_{h}\in V_{h}}\|\chi_{h}-w\|_{W^{1,p}(\Omega)},\qquad\forall w\in W_{0}^{1,p}(\Omega).$
@@
If the domain $\Omega$ is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), the estimate (30) holds for any $p\in(1,+\infty]$. If the domain is a Lipschitz polygon in $\mathbb{R}^{2}$, the estimate (30) holds for any $p\in(\frac{4}{3}-\epsilon_{1},+\infty]$ where $\epsilon_{1}$ is the same as that in Lemma 2.1. The constants $C$ and $h_{0}$ may depend $\tilde{A}_{0}$.
@@
By [25, Theorem $2$] ($d=3$), [38, Theorem $2$] ($d=2$),
@@
if $h<h_{0}$. Here the constant $C$ may depend on $\tilde{A}_{0}$. Therefore, there is a constant $C$ such that for any $2\leq p\leq+\infty$, if $h<h_{0}$,
@@
(32, 33) imply that for any $p\in(1,+\infty]$,
@@
We denote by $\boldsymbol{\phi}_{h}$ the standard $L^{2}$-orthogonal projection of $\nabla w$ into $P_{0}(\mathcal{T}_{h}]^{d}$. Then by the discrete trace inequality, we have
@@
Then by (37,38) and the discrete inverse inequality, we can conclude that the proof is complete. ∎
@@
By (28), $\mathcal{L}_{\tilde{A}_{0},h}w_{h}$ is the same as $\mathcal{L}_{0,h}$ in [14, Lemma $2.7$]. Therefore, by replacing [14, Lemma $2.6$] by Lemma 3.3, the proof in [14, Lemma $2.7$] is also valid for this Lemma. It is easy to verify that the proof of [14, Lemma $2.7$] is independent of the choice of $\boldsymbol{x}_{0}$, $R$ and $R^{\prime}$. ∎
@@ 3.2.2. Stability estimates for linear elliptic PDE in non-divergence form with uniformly continuous AA
We dived the proof into two steps.
@@
We notice that $I_{h}(\eta v_{h})=\eta v_{h}=0$ in $\Omega\setminus B_{\frac{11}{6}R}$ for any $v_{h}\in V_{h}$. Thus [14, Lemma $2.4$] implies that for any $q\in(1,+\infty)$, there is a constant $C_{q}$ such that
@@
Therefore, (46,48) imply that
@@
By [14, Lemma $2.2$], we have
$\|_{W_{h}^{2,p}(B_{\frac{11}{6}R})}\leq Ch^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
@@
By [14, Lemma $2.2$], we have
$\|_{W_{h}^{2,p}(B_{\frac{23}{12}R})}\leq Ch^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
@@
We notice $I_{h}(\eta v_{h})\in V_{h}(B_{\frac{11}{6}R})$. Then by (40,8,44,47), we have
@@
Since
$\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1},\qquad\forall h>0,$
there is $w\in W_{0}^{1,p}(\Omega)$ such that
@@
for a subsequence of $h\rightarrow 0$. We denote by $\boldsymbol{\phi}_{h}\in[H^{1}(\Omega)\cap\overline{V}_{h}]^{d}$ the standard $H^{1}$-conforming interpolation of $\nabla w_{h}$. By Lemma 2.3, we have
@@
(57) and the fact $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1}$ imply that there is $\boldsymbol{\phi}\in[W^{1,p}(\Omega)]^{d}$ such that
@@
for a subsequence of $h\rightarrow 0$. (56, 58) imply that $\nabla w=\boldsymbol{\phi}$ almost everywhere in $\Omega$. Therefore, we have that $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ and
@@
According to (59, 60, 62), We can conclude the proof is complete. ∎
@@
(63) $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)},\qquad\forall w_{h}\in V_{h}.$
@@
(73) $\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}\leq C\big(\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big).$
@@
We divide the proof into two steps. In step $1$, We will first prove the estimates (71) and (72,73) for $1<p\leq 2$ if $\Omega$ is convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), and $\frac{4}{3}-\epsilon_{1}<p<\frac{4}{3}+\epsilon_{2}$ if $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$. In step $2$, we will prove the estimate (71) for $1<p\leq\frac{4}{3}-\epsilon_{1}$ and $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$.
@@ 3.4. Linear elliptic PDE in non-divergence form with discontinuous $A$ but $\gamma A$ dominated by $I_{d}$
(77) $\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}\leq C\big(\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big),$
@@ 4.1. Auxiliary results for HJB equation
On the other hand, (80a,81a, 81b) implies
@@
The following Lemma 4.3 shows that the assumptions on coefficients and source term of (2) used in [40] and ([23, Proposition $4.3$] imply the assumption (6).
@@ 4.2. Well-posedness of Hamilton-Jacobi-Bellman equation
When the domain is convex and $p=2$, the condition (83) for coefficients in Theorem 4.1 is identical to the Cordès condition used in [40].
@@
for some $\epsilon\in[\underline{\kappa}^{\prime},1)$ where $\underline{\kappa}^{\prime}\in[0,1)$. In addition, $\underline{\kappa}^{\prime}=0$ if $p=2$ and the domain is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). The proof is similar. We would like to point out that the condition (83) and the condition (84) are exactly the Cordès conditions in [40].
@@
where the constant $\underline{C}_{p,\lambda}$ may depend on $p$ and $\lambda$. In fact, $\underline{C}_{2,\lambda}=1$ for any $\lambda\geq 1$ with convex domain due to [40, Theorem $2$].
@@
By (83) and the proof of [40, Lemma $1$] ($\underline{\kappa}$ will be determined later), we have that for any $w_{1},w_{2}\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
@@
In fact, $\underline{C}_{2,\lambda}^{\prime}=1$ with convex domain.
By (88,90), we have
@@
By (88) and (90), if (83) is satisfied with $\epsilon\in(\underline{\kappa},1)$, $T_{\lambda}$ is a contraction. Then there is a unique solution $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ satisfying the Hamilton-Jacobi-Bellman equation (2) which is equivalent to (87). Since $\underline{C}_{2,\lambda}=\underline{C}_{2,\lambda}^{\prime}=1$, $\underline{\kappa}=1$ if $p=2$ with convex domain.

### v3

@@ 3.1. Global $W^{2,p}$ estimate of linear elliptic PDE in non-divergence form with uniformly continuous AA
Since $\boldsymbol{x}_{0}\in\overline{\Omega}$ is chosen arbitrarily, we can conclude that $\phi\in W^{2,q}(\Omega)$ where $q=\min(\dfrac{3p}{3-p},2)>\frac{3}{2}$. So $\phi\in C^{0}(\overline{\Omega})$. Furthermore, by the interior estimate in [25, Lemma $9.16$], we have $\phi\in W_{\text{loc}}^{2,3}(\Omega)$ since $A:D^{2}\phi+\boldsymbol{b}\cdot\nabla\phi+c\phi=0$ in $\Omega$. Therefore, by the ABP estimate [25, Theorem $9.1$], we can conclude that $\phi=0$ in $\Omega$. ∎
@@
When $d=2$, $W^{2,p}(\Omega)\Subset C^{0}(\overline{\Omega})$ due to Sobolev’s embedding theorem. Furthermore, by the interior estimate in [25, Lemma $9.16$], we have $u\in W_{\text{loc}}^{2,2}(\Omega)$. Therefore, Theorem 3.1 can be proven by mimicking the proofs of [37, Theorem $1.1$] and [37, Corollary $1.2$] with the global $W^{2,p}$ estimate of Poisson equation (see [8, Proposition $2.1$]) and the Alexandroff–Bakelman–Pucci estimate estimate [25, Theorem $9.1$].
Now we focus on the case $d=3$. Due to [37, Corollary $1.2$], Theorem 3.1 holds if we further assume $\Omega$ is a convex polyhedra in $\mathbb{R}^{3}$ and $\frac{3}{2}<p\leq 2$. In fact, if we replace [11, Corollary $3.12$] by Lemma 2.1 in the proof of [37, Theorem $1.1$], it is easy to have that for any $1<p\leq 2$ and for any $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
@@
As the proof of [37, Corollary $1.2$], Theorem 3.1 will be proven via (33) and the proof of contraction with the uniqueness of solution in $W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$:
@@
When $\frac{3}{2}<p\leq 2$, it is straightforward by the Alexandroff–Bakelman–Pucci estimate estimate with the fact $W^{2,p}(\Omega)\Subset C^{0}(\overline{\Omega})$ (see the proof of [37, Corollary $1.2$]). Therefore, in order to complete the proof of Theorem 3.1, we only need the uniqueness for $1<p\leq\frac{3}{2}$, which is provided by Lemma 3.1. Therefore, we can conclude that the proof of Theorem 3.1 is complete. ∎
@@ 3.2. Stability analysis of FEM for linear elliptic PDE in non-divergence form with uniformly continuous AA
We define a linear operator $\mathcal{L}_{\tilde{A}_{0},h}:W_{h}^{1}\rightarrow V_{h}$ as
@@
(36) $\|w_{h}-w\|_{W^{1,p}(\Omega)}\leq C\inf_{\chi_{h}\in V_{h}}\|\chi_{h}-w\|_{W^{1,p}(\Omega)},\qquad\forall w\in W_{0}^{1,p}(\Omega).$
@@
If the domain $\Omega$ is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), the estimate (36) holds for any $p\in(1,+\infty]$. If the domain is a Lipschitz polygon in $\mathbb{R}^{2}$, the estimate (36) holds for any $p\in(\frac{4}{3}-\epsilon_{1},+\infty]$ where $\epsilon_{1}$ is the same as that in Lemma 2.1. The constants $C$ and $h_{0}$ may depend on $\tilde{A}_{0}$.
@@
Obviously, we have
(37) $\|\nabla(w_{h}-w)\|_{L^{2}(\Omega)}\leq C\|\nabla(\chi_{h}-w)\|_{L^{2}(\Omega)},\qquad\forall\chi_{h}\in V_{h}.$
By [26, Theorem $2$] ($d=3$), [41, Theorem $2$] ($d=2$),
@@
if $h<h_{0}$. Here the constant $C$ in (37,38) may depend on $\tilde{A}_{0}$. Therefore, according to (37,38), there is a constant $C$ such that for any $2\leq p\leq+\infty$, if $h<h_{0}$,
@@
(39, 40) imply that for any $p\in(1,+\infty]$,
@@
We denote by $\boldsymbol{\phi}_{h}$ the standard $L^{2}$-orthogonal projection of $\nabla w$ into $P_{0}(\mathcal{T}_{h})]^{d}$. Then by the discrete trace inequality, we have
@@
Then by (44,45) and the discrete inverse inequality, we can conclude that the proof is complete. ∎
@@
By (34), $\mathcal{L}_{\tilde{A}_{0},h}w_{h}$ is the same as $\mathcal{L}_{0,h}$ in [15, Lemma $2.7$]. Therefore, by replacing [15, Lemma $2.6$] by Lemma 3.3, the proof in [15, Lemma $2.7$] is also valid for this Lemma. It is easy to verify that the proof of [15, Lemma $2.7$] is independent of the choice of $\boldsymbol{x}_{0}$, $R$ and $R^{\prime}$. ∎
@@ 3.2.2. Stability estimates for linear elliptic PDE in non-divergence form with uniformly continuous AA
We divided the proof into two steps.
@@
We notice that $I_{h}(\eta v_{h})=\eta v_{h}=0$ in $\Omega\setminus B_{\frac{11}{6}R}$ for any $v_{h}\in V_{h}$. Thus [15, Lemma $2.4$] implies that for any $q\in(1,+\infty)$, there is a constant $C_{q}$ such that
@@
Therefore, (53,55) imply that
@@
By [15, Lemma $2.2$], we have
$\|w_{h}\|_{W_{h}^{2,p}(B_{\frac{11}{6}R})}\leq Ch^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
@@
By [15, Lemma $2.2$], we have
$\|w_{h}\|_{W_{h}^{2,p}(B_{\frac{23}{12}R})}\leq Ch^{-1}\|w_{h}\|_{W^{1,p}(B_{2R})}.$
@@
We notice $I_{h}(\eta v_{h})\in V_{h}(B_{\frac{11}{6}R})$. Then by (47,8,51,54), we have
@@
Since $\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq M_{1}$ for any $h>0$, Lemma 2.4 implies
$\|w_{h}\|_{W^{1,p}(\Omega)}\leq CM_{1},\qquad\forall h>0.$
Thus there is $w\in W_{0}^{1,p}(\Omega)$ such that
@@
for a subsequence of $h\rightarrow 0$. We denote by $\boldsymbol{\phi}_{h}\in[H^{1}(\Omega)\cap\overline{V}_{h}]^{d}$ the standard $H^{1}$-conforming averaging of $\nabla w_{h}$. By Lemma 2.3, we have
@@
(64) and the fact $\|w_{h}\|_{W^{1,p}(\Omega)}+\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq(1+C)M_{1}$ imply that there is $\boldsymbol{\phi}\in[W^{1,p}(\Omega)]^{d}$ such that
@@
for a subsequence of $h\rightarrow 0$. (63,65) imply that $\nabla w=\boldsymbol{\phi}$ almost everywhere in $\Omega$. Therefore, we have that $w\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ and
@@
According to (66, 67, 69), We can conclude the proof is complete. ∎
@@
(70) $\|w_{h}\|_{W^{1,p}(\Omega)}+\|w_{h}\|_{W_{h}^{2,p}(\Omega)}\leq C\|\tilde{A}:\overline{\nabla}_{h}(\nabla w_{h})+\tilde{\boldsymbol{b}}\cdot\nabla w_{h}+\tilde{c}w_{h}\|_{L_{h}^{p}(\Omega)},\quad\forall w_{h}\in V_{h}.$
@@
(80) $\|u_{h}-u\|_{W^{1,p}(\Omega)}+\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}$
$\leq C$ $\big(\|u-\chi_{h}\|_{W^{1,p}(\Omega)}+\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big).$
@@
We divide the proof into two steps. In step $1$, We will first prove the estimates (78) and (79,80) for $1<p\leq 2$ if $\Omega$ is convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$), and $\frac{4}{3}-\epsilon_{1}<p<\frac{4}{3}+\epsilon_{2}$ if $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$. In step $2$, we will prove the estimate (78) for $1<p\leq\frac{4}{3}-\epsilon_{1}$ and $\Omega$ is a polygon (maybe non-convex) in $\mathbb{R}^{2}$.
@@ 3.4. Linear elliptic PDE in non-divergence form with discontinuous $A$ but $\gamma A$ dominated by $I_{d}$
(84) $\|u_{h}-u\|_{W^{1,p}(\Omega)}+\|u_{h}-u\|_{W_{h}^{2,p}(\Omega)}$
$\leq$ $C\big(\|u-\chi_{h}\|_{W^{1,p}(\Omega)}+\|u-\chi_{h}\|_{W_{h}^{2,p}(\Omega)}+\|\overline{P}_{h}(D^{2}u)-D^{2}u\|_{L^{p}(\Omega)}\big),$
@@ 4.1. Auxiliary results for HJB equation
On the other hand, (87a,88a, 88b) implies
@@
The following Lemma 4.3 shows that the assumptions on coefficients and source term of (2) used in [43] and ([24, Proposition $4.3$] imply the assumption (6).
@@ 4.2. Well-posedness of Hamilton-Jacobi-Bellman equation
When the domain is convex and $p=2$, the condition (90) for coefficients in Theorem 4.1 is identical to the Cordes condition used in [43].
@@
for some $\epsilon\in[\underline{\kappa}^{\prime},1)$ where $\underline{\kappa}^{\prime}\in[0,1)$. In addition, $\underline{\kappa}^{\prime}=0$ if $p=2$ and the domain is a convex polyhedra in $\mathbb{R}^{d}$ ($d=2,3$). The proof is similar. We would like to point out that the condition (90) and the condition (91) are exactly the Cordes conditions in [43].
@@
where the constant $\underline{C}_{p,\lambda}$ may depend on $p$ and $\lambda$. In fact, $\underline{C}_{2,\lambda}=1$ for any $\lambda\geq 1$ with convex domain due to [43, Theorem $2$].
@@
By (90) and the proof of [43, Lemma $1$] ($\underline{\kappa}$ will be determined later), we have that for any $w_{1},w_{2}\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$,
@@
In fact, $\underline{C}_{2,\lambda}^{\prime}=1$ on convex domain.
By (95,97), we have
@@
By (95) and (97), if (90) is satisfied with $\epsilon\in(\underline{\kappa},1)$, $T_{\lambda}$ is a contraction. Then there is a unique solution $u\in W^{2,p}(\Omega)\cap W_{0}^{1,p}(\Omega)$ satisfying the Hamilton-Jacobi-Bellman equation (2) which is equivalent to (94). Since $\underline{C}_{2,\lambda}=\underline{C}_{2,\lambda}^{\prime}=1$, $\underline{\kappa}=0$ if $p=2$ with convex domain.